# Heights and periods of algebraic cycles in families

### v1

@@ 6.2. Relating with the Gross–Schoen normal function
By (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ Hence, they define the same normal functions $\nu_{{\mathrm{GS}}}$ on $U$. Set
@@ 6.3. Bigness of the generic fiber of the adelic line bundle
By [YZ21, Thm. 5.3.5], lower bounds of the height function $h_{\overline{{\mathscr{L}}}}$ correspond to bigness properties of $\overline{{\mathscr{L}}}$. In this paper, we prove the bigness of the generic fiber of $\overline{{\mathscr{L}}}$ and deduce the desired height comparison from it.
There is a natural base change of adelic line bundles $\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}\rightarrow\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Q}})_{{\mathbb{Q}}}$. Denote by $\tilde{{\mathscr{L}}}$ the image of $\overline{{\mathscr{L}}}$ under this base change map; $\tilde{{\mathscr{L}}}$ is called the generic fiber of $\overline{{\mathscr{L}}}$.
We start by stating the following formula for $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})$. Notice that it does not follow from arithmetic Hilbert–Samuel because we have not proved the nefness of $\tilde{{\mathscr{L}}}$.
Proposition 6.3.
We have
$\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})^{3g-3}.$
The proof of this proposition shall be postponed to §$\S$6.4.
@@
Assume $g\geq 3$. Then the adelic line bundle $\tilde{{\mathscr{L}}}$ is big, i.e. $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})>0$.
@@
By Proposition 6.3 and Proposition 6.2, we have $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}$. Now $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\geq 0$ since $\beta_{{\mathrm{GS}}}$ is semi-positive. By Corollary 5.3, $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\not\equiv 0$. Hence $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})>0$. ∎
@@
The claim for $\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}$ is a direct consequence of Theorem 6.4 and [YZ21, Thm. 5.3.5.(iii)]. Notice that one can either take the adelic line bundle on ${\mathscr{M}}_{g}$, which defines the Faltings height (it exists by [YZ21, §$\S$2.6.2]), or one can take any ample line bundle on a suitable compactification of ${\mathscr{M}}_{g}$ and then use the comparison of the logarithmic Weil height and the Faltings height.
@@
The claim for $\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}$ follows because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple. ∎
6.4. Proof of Proposition 6.3
Denote for simplicity by ${\mathscr{S}}={\mathscr{M}}_{g}$.
@@
By the flatness of extension ${\mathbb{Q}}\subseteq{\mathbb{C}}$, we have $\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{\mathbb{Q}})=\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{\mathbb{C}})$. By [YZ21, Theorem 5.2.1], it suffices to construct a sequence of model Hermitian line bundles $({\mathscr{S}}_{i},\overline{\mathscr{L}}_{i})$ with limit $({\mathscr{S}},\overline{\mathscr{L}})$ so that
$\lim_{i{\longrightarrow}\infty}\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{i})=\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})^{3g-3}.$
Let $\overline{\mathscr{N}}$ be an ample Hermitian line bundle on some integral model ${\mathscr{S}}$ of ${\mathscr{S}}$. Following proof of [YZ21, Theorem 6.1.1], we have a sequence of model line bundles $({\mathscr{J}}_{i},\overline{\mathscr{P}}^{\Delta}_{i},\ell_{i})$ with the limit $({\mathscr{J}},\overline{\mathscr{P}}^{\Delta})$, so that $\overline{\mathscr{P}}^{\Delta}_{i}+4^{-i}\pi^{*}\overline{\mathscr{N}}$ is nef.
@@
For any $n\in{\mathbb{N}}$, we have an action of ${\mathscr{J}}[n]$ on ${\mathscr{J}}$ by translating torsion points:
$m_{n}\colon{\mathscr{J}}[n]\times{\mathscr{J}}\overset{(m,p)}{{\longrightarrow}}{\mathscr{J}}\times{\mathscr{J}},\qquad(t,x)\mapsto(x+t,x)$
@@
By the Theorem of square, ${\mathrm{N}}_{p}m_{n}^{*}{\mathscr{P}}^{\Delta}=({\mathscr{P}}^{\Delta})^{n^{2g}}$. Thus, we obtain an adelic metrized line bundle
$\overline{\mathscr{P}}^{\Delta}_{i,n}:=n^{-2g}{\mathrm{N}}_{p}m_{n}^{*}\overline{\mathscr{P}}^{\Delta}_{i}.$
@@
This bundle is again realized on some model ${\mathscr{J}}_{i,n}$ of ${\mathscr{J}}$ with connection morphism $\ell_{i,n}:{\mathscr{P}}^{\Delta}{\longrightarrow}{\mathscr{P}}^{\Delta}_{i,n,{\mathscr{J}}}$. Moreover ${\mathrm{div}}(\ell_{i,n})$ is in fact bounded by ${\mathrm{div}}(\ell_{i})$. Over ${\mathbb{C}}$, the curvature form $c_{1}(\overline{\mathscr{P}}^{\Delta}_{i,n,{\mathbb{C}}})$ is obtained from the curvature form $c_{1}(\overline{\mathscr{P}}^{\Delta}_{i,{\mathbb{C}}})$ by taking over average over $n$-torsion points. It follows that these forms converge to $c_{1}(\overline{\mathscr{P}}^{\Delta})$ uniformly in any compact subset of ${\mathscr{J}}({\mathbb{C}})$. These bundles also induce a double sequence of model line bundles $({\mathscr{S}}_{i},\overline{\mathscr{L}}_{i,n})$ of $({\mathscr{S}},\overline{\mathscr{L}})$ so that they convergent to $({\mathscr{S}},\overline{\mathscr{L}})$ as $i{\longrightarrow}\infty$, and that the metric $c_{1}(\overline{\mathscr{L}}_{i,n,{\mathbb{C}}})$ uniformly convergent to $c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})$ as $n{\longrightarrow}\infty$.
@@
More precisely, we let $\Omega_{i}$ be an increasing sequence of relatively compact open subsets of ${\mathscr{S}}({\mathbb{C}})$ so that ${\mathscr{S}}({\mathbb{C}})=\cup\Omega_{i}$ and $\epsilon_{i}$ be a decreasing sequence of positive numbers convergent to 00 so that on $\Omega_{i}$,
$c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})\leq\epsilon_{i}^{-1}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
Then for each $i$, we choose $n_{i}$ so that
@@
as hermitian forms on the tangent bundle on $\Omega_{i}$. We simply write ¯$\overline{\mathscr{L}}_{i}=\overline{\mathscr{L}}_{i,n_{i}}$. Using the reference curvature $c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})$, we may talk about eigenvalues of $c_{1}(\overline{\mathscr{L}}_{{\mathbb{C}}})$ and $c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})$.
Now we apply Demailly’s Morse inequality [Dem91] for the bundle ¯$\overline{\mathscr{L}}_{i}$ on ${\mathscr{S}}_{i}({\mathbb{C}})$. For each $q\in{\mathbb{N}}$, let ${\mathscr{S}}_{i,q}$ denote the subset of ${\mathscr{S}}_{i}({\mathbb{C}})$ of points where $c_{1}(\overline{\mathscr{L}}_{i})$ has $q$-negative eigenvalues and $n-q$ positive eigenvalues. Then by [Dem91, (1.3), (1.5)], we have the following estimate as $k{\longrightarrow}\infty$:
$h^{q}(k{\mathscr{L}}_{i})\leq\frac{k^{d}}{d!}\left|\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}\right|+o(k^{d}),\qquad d:=3g-3$
$\sum_{q}(-1)^{q}h^{q}(k{\mathscr{L}}_{i})=\frac{k^{d}}{d!}\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}+o(k^{d}).$
@@
$\left|\widehat{\mathrm{vol}}({\mathscr{L}}_{i})-\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}\right|\leq\sum_{q>0}\left|\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}\right|.$
By [YZ21, Theorem 5.4.4],
$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}=\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}})^{d}.$
Thus, it remains to the following estimate for each $q>0$:
$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}=0.$
We want to prove this on $\Omega_{i}\cap{\mathscr{S}}_{i,q}$ and on its complement ${\mathscr{S}}_{i,q}\setminus\Omega_{i}$ respectively.
By (6.4), we have
$-\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\leq c_{1}(\overline{\mathscr{L}}_{i})\leq(\epsilon^{-1}+\epsilon^{d})c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
Thus on $\Omega_{i}\cap{\mathscr{S}}_{i,q}$, $c_{1}(\overline{\mathscr{L}}_{i})$ has all eigenvalue $\leq\epsilon^{-1}+\epsilon^{d}$ and one negative eigenvalue with absolute value bounded by $\leq\epsilon^{d}$. It follows that $|c_{1}({\mathscr{L}}_{i})^{d}|$ is bounded by
$\epsilon^{d}(\epsilon^{-1}+\epsilon^{d})^{d-1}c_{1}(\overline{\mathscr{N}})^{d}=\epsilon_{i}(1+\epsilon_{i}^{d+1})^{d-1}c_{1}(\overline{\mathscr{N}})^{d}.$
It follows that
$\int_{{\mathscr{S}}_{i,q}\cap\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{d}|=O(\epsilon_{i}).$
It remains to treat the integral over ${\mathscr{S}}_{i,q}\setminus\Omega_{i}$. Using decomposition

### v2

@@ 6.2. Relating with the Gross–Schoen normal function
By (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$ Hence, they define the same normal function $\nu_{{\mathrm{GS}}}$ on $U$. Set
@@ 6.3. Bigness of the generic fiber of the adelic line bundle
By [YZ21, Thm. 5.3.5] (which we reformulate as Theorem E.9), lower bounds of the height function $h_{\overline{{\mathscr{L}}}}$ correspond to bigness properties of $\overline{{\mathscr{L}}}$. In this paper, we prove the bigness of the generic fiber $\widetilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$ and deduce the desired height comparison from it. The definition of $\widetilde{{\mathscr{L}}}$ is recalled in §$\S$§E.5. Roughly speaking, there is a natural base change of adelic line bundles $\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}\rightarrow\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Q}})_{{\mathbb{Q}}}$, and $\widetilde{{\mathscr{L}}}$ is the image of $\overline{{\mathscr{L}}}$ under this base change.
The arithmetic volume of $\widetilde{{\mathscr{L}}}$ is defined in (E.11). We will prove the following formula for $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}})$. Notice that it does not follow from arithmetic Hilbert–Samuel because we have yet to prove the nefness of $\widetilde{{\mathscr{L}}}$.
@@
Let $S\subseteq{\mathscr{M}}_{g}$ be an irreducible subvariety, and let ~$\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S}$ be the image of $\widetilde{{\mathscr{L}}}$ under the natural base change $\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Q}})_{{\mathbb{Q}}}\rightarrow\widehat{\mathrm{Pic}}(S/{\mathbb{C}})_{{\mathbb{Q}}}$. Then we have
(6.4) $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S})=\int_{S({\mathbb{C}})}c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})^{\wedge\dim S}=\int_{S({\mathbb{C}})}\beta_{{\mathrm{GS}}}^{\wedge\dim S}.$
The proof of this theorem shall be postponed to §$\S$6.4. An interesting result of this volume computation is as follows. For each subvariety $S$ of ${\mathscr{M}}_{g}$, consider the Betti strata (1.2) with $\nu=\nu_{{\mathrm{GS}}}$.
Corollary 6.4.
Let $S$ be a subvariety of ${\mathscr{M}}_{g}$ defined over a subfield $k$ of ${\mathbb{C}}$. Then $S^{\mathrm{Betti}}(1)$ is also defined over $k$. In particular ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is defined over ${\mathbb{Q}}$.
@@
Let $S^{\prime}$ be an irreducible component of $S^{\mathrm{Betti}}(1)$. Then $S^{\prime}=S^{\prime}_{{\mathscr{F}}}(1)$, and hence $\beta_{{\mathrm{GS}}}^{\wedge\dim S^{\prime}}\equiv 0$ on $S^{\prime}$; see (1.6) and Corollary D.7. So $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S^{\prime}})=0$ by Theorem 6.3. By Definition E.11, for any automorphism $\sigma\in{\mathrm{Aut}}({\mathbb{C}}/k)$ we have
$\widehat{\mathrm{vol}}(\widetilde{\mathscr{L}}_{{\mathbb{C}}}\otimes_{\sigma}{\mathbb{C}}|_{S^{\prime}\otimes_{\sigma}{\mathbb{C}}})=\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S^{\prime}})=0.$
Let $S^{\prime\sigma}$ denote $S^{\prime}\otimes_{\sigma}{\mathbb{C}}$. Then $S^{\prime\sigma}$ is also a closed subvariety of $S$. As $\widetilde{\mathscr{L}}$ is defined over ${\mathbb{Q}}$, Proposition 6.3 implies
$\int_{S^{\prime\sigma}({\mathbb{C}})}c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})^{\wedge\dim S^{\prime}}=0.$
Thus $\beta_{{\mathrm{GS}}}^{\wedge\dim S^{\prime}}\equiv 0$ on $S^{\prime\sigma}$ by Proposition 6.2, and therefore
$S^{\prime\sigma}({\mathbb{C}})=S^{\prime\sigma}_{\mathscr{F}}(1)\subseteq S^{\mathrm{Betti}}(1).$
This shows that $S^{\mathrm{Betti}}(1)\otimes_{\sigma}{\mathbb{C}}=S^{\mathrm{Betti}}(1)$. Thus $S^{\mathrm{Betti}}(1)$ is defined over $k$. ∎
Theorem 6.5.
Assume $g\geq 3$. Then the adelic line bundle $\widetilde{{\mathscr{L}}}$ is big, i.e. $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}})>0$.
Proof.
By the flatness of extension ${\mathbb{Q}}\subseteq{\mathbb{C}}$, we have $\widehat{\mathrm{vol}}(\widetilde{\mathscr{L}})=\widehat{\mathrm{vol}}(\widetilde{\mathscr{L}}_{\mathbb{C}})$. So $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}$ by Theorem 6.3. Now $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\geq 0$ since $\beta_{{\mathrm{GS}}}$ is semi-positive, and $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\not\equiv 0$ by Corollary 5.3. So $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}})>0$. ∎
Now, we are ready to prove our main theorem.
@@
It suffices to prove the claim for $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}$ because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple.
Next by [Zha10, Cor. 2.5.2], it suffices to prove for $\xi=\xi_{s}$. We start by showing there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$ such that
(6.5) $\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{{\mathrm{Fal}}}(s)-c$
for all $s\in U(\overline{\mathbb{Q}})$. The conclusion follows immediately from Theorem 6.5 and [YZ21, Thm. 5.3.5.(iii)] (which we reformulated as Theorem E.9). Notice that one can either take the adelic line bundle on ${\mathscr{M}}_{g}$, which defines the Faltings height (it exists by [YZ21, §$\S$2.6.2]), or one can take any ample line bundle on a suitable compactification of ${\mathscr{M}}_{g}$ and then use the comparison of the logarithmic Weil height and the Faltings height.
We can do better to get an explicit $U$. Suppose $U_{1},U_{2}$ are open subsets of ${\mathscr{M}}_{g}$ over ${\mathbb{Q}}$ such that (6.5) holds true for all $s\in U_{1}(\overline{\mathbb{Q}})$ and $s\in U_{2}(\overline{\mathbb{Q}})$, then (6.5) holds true for all $s\in(U_{1}\cup U_{2})(\overline{\mathbb{Q}})$. Therefore, there exists a maximal Zariski open dense subset $U^{\prime}$ of ${\mathscr{M}}_{g}$ on which (6.5) holds. Let $S:={\mathscr{M}}_{g}\setminus U^{\prime}$. Then $S$ is proper Zariski closed in ${\mathscr{M}}_{g}$, and the height inequality (6.5) cannot hold for any non-empty open subvariety $V$ of $S$. By Theorem E.9, $\widetilde{{\mathscr{L}}}$ cannot be big on any irreducible component $Z$ of $S$, and hence $c_{1}(\overline{\mathscr{L}}|_{Z({\mathbb{C}})})^{\wedge\dim Z}\equiv 0$ by Proposition 6.3. Thus $Z_{{\mathbb{C}}}=Z_{{\mathscr{F}}}(1)\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. So $S_{{\mathbb{C}}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. So $S\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ by Corollary 6.4. Thus we can take ${\mathscr{M}}_{g}^{\mathrm{amp}}$ to be ${\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, which is Zariski open dense and is defined over ${\mathbb{Q}}$, in Theorem 1.1.∎
6.4. Proof of Theorem 6.3
The second equality is immediate by Proposition 6.2. We will prove the first equality. We start to work on ${\mathscr{M}}_{g}$. To ease notation, denote by ${\mathscr{M}}={\mathscr{M}}_{g}$ and ${\mathscr{J}}={\mathscr{J}}_{g}$.
@@
By [YZ21, Theorem 5.2.1], it suffices to construct a sequence of model (Hermitian) line bundles $({\mathscr{M}}_{i},\overline{\mathscr{L}}_{i})$ which converges to $({\mathscr{M}},\overline{\mathscr{L}})$ (see the beginning of §$\S$§E.1 for notation) such that
(6.6) $\lim_{i\rightarrow\infty}\mathrm{vol}({\mathscr{L}}_{i,{\mathbb{C}}}|_{S_{i}})=\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})^{\wedge\dim S},$
where $S_{i}$ is the Zariski closure of $S\subseteq{\mathscr{M}}\subseteq{\mathscr{M}}_{i,{\mathbb{C}}}$.
We start the construction. First, the proof of [YZ21, Thm. 6.1.1] yields
  • a sequence of projective integral models $\pi_{i}\colon{\mathscr{J}}_{i}\rightarrow{\mathscr{M}}$ of ${\mathscr{J}}\rightarrow{\mathscr{M}}$,
  • an ample Hermitian line bundle $\overline{\mathscr{N}}$ on ${\mathscr{M}}$,
  • a sequence of model line bundles $({\mathscr{J}}_{i},\overline{\mathscr{P}}^{\Delta}_{i})$ converging to $({\mathscr{J}},\overline{\mathscr{P}}^{\Delta})$,
such that $\overline{\mathscr{P}}^{\Delta}_{i}+4^{-i}\pi_{i}^{*}\overline{\mathscr{N}}$ is nef. Denote by $\ell_{i}$ the connection morphism ${\mathscr{P}}^{\Delta}\xrightarrow{\sim}{\mathscr{P}}^{\Delta}_{i,{\mathscr{J}}}:={\mathscr{P}}^{\Delta}_{i}\times_{{\mathscr{J}}_{i}}{\mathscr{J}}$.
@@
For each $n\in{\mathbb{N}}$, we have an action of ${\mathscr{J}}[n]$ on ${\mathscr{J}}$ by translating torsion points:
$m_{n}\colon{\mathscr{J}}[n]\times_{S}{\mathscr{J}}{\longrightarrow}{\mathscr{J}},\qquad(t,x)\mapsto x+t.$
@@
By Theorem of the square, the norm line bundle ${\mathrm{N}}_{p}m_{n}^{*}{\mathscr{P}}^{\Delta}$ is $({\mathscr{P}}^{\Delta})^{\otimes n^{2g}}$, where $p$ is the natural projection ${\mathscr{J}}[n]\times_{S}{\mathscr{J}}\rightarrow{\mathscr{J}}$. Thus, we obtain an adelic line bundle
$\overline{\mathscr{P}}^{\Delta}_{i,n}:=n^{-2g}{\mathrm{N}}_{p}m_{n}^{*}\overline{\mathscr{P}}^{\Delta}_{i}$
@@
which is realized on some model ${\mathscr{J}}_{i,n}$ of ${\mathscr{J}}$ with connection morphism $\ell_{i,n}:{\mathscr{P}}^{\Delta}\xrightarrow{\sim}{\mathscr{P}}^{\Delta}_{i,n,{\mathscr{J}}}$. Moreover ${\mathrm{div}}(\ell_{i,n})$ is in fact bounded by ${\mathrm{div}}(\ell_{i})$. Over ${\mathbb{C}}$, the curvature form $c_{1}(\overline{\mathscr{P}}^{\Delta}_{i,n,{\mathbb{C}}})$ is obtained from the curvature form $c_{1}(\overline{\mathscr{P}}^{\Delta}_{i,{\mathbb{C}}})$ by taking over average over $n$-torsion points. It follows that these forms converge to $c_{1}(\overline{\mathscr{P}}^{\Delta})$ uniformly in any compact subset of ${\mathscr{J}}({\mathbb{C}})$. These bundles also induce a double sequence of model line bundles $({\mathscr{M}}_{i},\overline{\mathscr{L}}_{i,n})$ of $({\mathscr{M}},\overline{\mathscr{L}})$ so that they convergent to $({\mathscr{M}},\overline{\mathscr{L}})$ as $i\rightarrow\infty$, and that the metric $c_{1}(\overline{\mathscr{L}}_{i,n,{\mathbb{C}}})$ convergent uniformly to $c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})$ as $n\rightarrow\infty$.
@@
More precisely, we let $\Omega_{i}$ be an increasing sequence of relatively compact open subsets of ${\mathscr{M}}({\mathbb{C}})$ with ${\mathscr{M}}({\mathbb{C}})=\cup\Omega_{i}$ and $\epsilon_{i}$ be a decreasing sequence of positive numbers convergent to 00 such that
(6.7) $c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})\leq\epsilon_{i}^{-1}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\quad\text{ on }\Omega_{i}.$
Then for each $i$, choose $n_{i}$ such that
@@
as Hermitian forms on the tangent bundle of $\Omega_{i}$.
For each $i\geq 1$, set ¯$\overline{\mathscr{L}}_{i}:=\overline{\mathscr{L}}_{i,n_{i}}$. We will show that the sequence $({\mathscr{M}}_{i},\overline{\mathscr{L}}_{i})$ is the desired sequence of model line bundles such that (6.6) holds.
Using the reference curvature $c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})$, we may talk about eigenvalues of $c_{1}(\overline{\mathscr{L}}_{{\mathbb{C}}})$ and $c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})$. Recall that $S_{i}$ is the closure of $S\subseteq{\mathscr{M}}\subseteq{\mathscr{M}}_{i,{\mathbb{C}}}$ for each $i$.
From now on, to ease notation we use ${\mathscr{L}}_{i,{\mathbb{C}}}$ to denote ${\mathscr{L}}_{i,{\mathbb{C}}}|_{S_{i}}$.
Let us apply Demailly’s Morse inequality [Dem91] to the bundle ¯$\overline{\mathscr{L}}_{i,{\mathbb{C}}}$ on $S_{i}({\mathbb{C}})$. For each $q\in{\mathbb{N}}$, let $S_{i,q}$ denote the subset of $S_{i}({\mathbb{C}})$ of points where $c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})$ has $q$ negative eigenvalues and $n-q$ positive eigenvalues. Then by [Dem91, (1.3), (1.5)], we have the following estimates as $k\rightarrow\infty$, with $d:=\dim S$
$h^{q}(k{\mathscr{L}}_{i,{\mathbb{C}}})\leq\frac{k^{d}}{d!}\left|\int_{S_{i,q}}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}\right|+o(k^{d}),$
$\sum_{q}(-1)^{q}h^{q}(k{\mathscr{L}}_{i,{\mathbb{C}}})=\frac{k^{d}}{d!}\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}+o(k^{d}).$
It follows, combined with the definition of $\mathrm{vol}({\mathscr{L}}_{i,{\mathbb{C}}})$, that
$\left|\mathrm{vol}({\mathscr{L}}_{i,{\mathbb{C}}})-\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}\right|\leq\sum_{q>0}\left|\int_{S_{i,q}}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}\right|.$
By [YZ21, Thm. 5.4.4],
$\lim_{i\rightarrow\infty}\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}=\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{{\mathbb{C}}})^{\wedge d}.$
Thus to prove (6.6), it remains to prove
(6.9) $\lim_{i\rightarrow\infty}\int_{S_{i,q}}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}=0\quad\text{ for each }q>0.$
We will prove (6.9) on $\Omega_{i}\cap S_{i,q}$ and on its complement $S_{i,q}\setminus\Omega_{i}$ respectively.
On $\Omega_{i}\cap S_{i,q}$, by (6.7) and (6.8) we have
$-\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\leq c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})\leq(\epsilon_{i}^{-1}+\epsilon_{i}^{d})c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
Thus on $\Omega_{i}\cap S_{i,q}$, $c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})$ has all eigenvalue $\leq\epsilon_{i}^{-1}+\epsilon_{i}^{d}$ and one negative eigenvalue with absolute value bounded by $\leq\epsilon_{i}^{d}$. It follows that $|c_{1}({\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}|$ is bounded by
$\epsilon_{i}^{d}(\epsilon_{i}^{-1}+\epsilon_{i}^{d})^{d-1}c_{1}(\overline{\mathscr{N}}_{{\mathbb{C}}})^{\wedge d}=\epsilon_{i}(1+\epsilon_{i}^{d+1})^{d-1}c_{1}(\overline{\mathscr{N}}_{{\mathbb{C}}})^{d}.$
@@
(6.10) $\int_{S_{i,q}\cap\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{\wedge d}|=O(\epsilon_{i}).$
On $S_{i,q}\setminus\Omega_{i}$, using